# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ 2 Setting and Notations
$\delta\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\;\delta\bm{\mathbf{\hat{W}}}^{t}-{\underline{\bm{G}}}\cdot\delta\bm{\Omega}^{t-1},$ (19)
$\delta\bm{\mathbf{\hat{W}}}^{t+1}$ $=\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\delta\bm{\Omega}^{t}.$ (20)
Substituting eq. (20) into eq. (20), this is equivalent to $\delta\bm{\Omega}^{t+1}={\underline{\bm{L}}}\cdot\delta\bm{\Omega}^{t}$. Moreover, assuming convergence, [^1] Here and in the rest of the paper, we drop the time index $t$ to refer to the quantities at convergence. one ends up with
$\delta\bm{\Omega}$ $={\underline{\bm{L}}}\cdot\delta\bm{\Omega}$ (21)
$\delta\bm{\mathbf{\hat{W}}}$ $={\underline{\bm{T}}}\cdot\delta\bm{\mathbf{\hat{W}}}.$ (22)
This suggests that at first leading order in the estimates, the dynamics is governed by power-iteration (and respectively a variant of it) on the tensor ${\underline{\bm{L}}}$ (respectively the symmetric tensor ${\underline{\bm{T}}}$) defined previously. As power iteration converges under mild assumptions to the top (matrix) eigenvector of the tensor, this further suggests to use the top eigenvectors $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}$, $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}$ of the corresponding tensors as an estimate for the weight matrix $\bm{W}_{\star}$.
Additional details on the linearized GAMP are reported in Appendix A.1. Similar approaches have been thoroughly investigated in the context of single-index models Mondelli and Montanari (2018); Maillard et al. (2022), community detection Krzakala et al. (2013); Saade et al. (2014), spiked matrix estimation Lesieur et al. (2017); Aubin et al. (2019), where they have been provably shown to provide a non-vanishing correlation with the ground truth exactly at the optimal weak recovery threshold. It is interesting to notice that the spectral estimators $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}$ and $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}$ correspond to the generalization for multi-index models of the spectral methods derived in Maillard et al. (2022), respectively from the linearization of the optimal Vector Approximate Message Passing Schniter et al. (2016); Rangan et al. (2017) and the Hessian of the TAP free energy associated to the posterior distribution for the weights Saade et al. (2014). In particular, for $p\!=\!1$, the matrices proposed in our manuscript exactly reduce to the two ones (called respectively "TAP" and "LAMP") investigated in Maillard et al. (2022).
@@ 3 Main Technical Results
In order to characterize the weak recovery of the proposed spectral methods, we define two message passing schemes tailored to respectively have eigenvectors of ${\underline{\bm{L}}}$ and ${\underline{\bm{T}}}$ as fixed points. Similarly to these previous works, we leverage the state evolution associated to the algorithms in order to quantify the alignment between the spectral estimators and the weight matrix $\bm{W}_{\star}$, tracking the overlap matrices
@@ 3.1 Asymmetric spectral method
Consider the linear GAMP algorithm (55,56) with denoiser functions
${\bm{g}}_{\rm out}^{t}(\bm{\omega},y)={\bm{G}}(y)\bm{\omega}\coloneqq\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}}\right)\bm{\omega},\quad\bm{f}_{in}^{t}(\bm{b})=\gamma^{-1}\bm{b},$ (24)
with $\gamma>0$ a parameter to be fixed. Its iterations read
$\bm{\Omega}^{t}=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-\gamma^{-1}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t-1},$ (25)
$\bm{\mathbf{\hat{W}}}^{t+1}=\gamma^{-1}\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t},$ (26)
Figure 1: Distribution of the eigenvalues (dots) $\lambda\in\mathbb{C}$ of ${\underline{\bm{L}}}$ at finite $n=10^{4}$, for $g(z_{1},z_{2})=z_{1}z_{2}$, $\alpha_{c}\approx 0.59375$. (Left) $\alpha=0.4<\alpha_{c}$. (Right) $\alpha=1>\alpha_{c}$. The dashed blue circle has radius equal to $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$, i.e. the value $\gamma_{b}$ predicted in Lemma 3.4. The dashed orange vertical line corresponds to $\operatorname{Re}\lambda=\nicefrac{{\alpha}}{{\alpha_{c}}}$, the eigenvalue $\gamma_{s}$ defined in Lemma 3.3. As predicted by the state evolution equations for this problem, two significant eigenvalues (highlighted in orange) are observed near this vertical line.
When $\gamma$ is chosen as an eigenvalue of ${\underline{\bm{L}}}$, the correspondent eigenvector is a fixed point of eq. (25). In the high-dimensional limit, the asymptotic overlap of the estimator $\bm{\mathbf{\hat{W}}}$ in (26) can be tracked thanks to the state evolution equations, which follow from an immediate application of the general result of Javanmard and Montanari (2013) to the GAMP algorithm (3.6).
Proposition 3.2 (State evolution Javanmard and Montanari (2013)).
@@
$\bm{M}^{t+1}$ $=\alpha\gamma^{-1}{\mathcal{F}}(\bm{M}^{t})$ (27)
@@
${\mathcal{F}}(\bm{M})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})],$ (29)
${\mathcal{G}}(\bm{M})$ $\coloneqq\mathbb{E}_{{\mathrm{y}}}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})\bm{M}^{T}{\bm{G}}({\mathrm{y}})].$ (30)
With the state evolution equations in hand, one can derive a sharp characterization of the asymptotic weak recovery threshold in terms of the spectral properties of the estimator by a linear stability argument Strogatz (2001)
Lemma 3.3.
For $\alpha>\alpha_{c}$, $\gamma_{s}=\nicefrac{{\alpha}}{{\alpha_{c}}}$ is the largest value of $\gamma$ such that the state evolution 3.2 has a stable fixed point ($\bm{M},\bm{Q}$) with
$\bm{M}\neq{\bm{0}},\quad\bm{Q}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}.$ (31)
Additionally, $\bm{M}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$.
Lemma 3.4.
For all $\alpha\in\mathbb{R}$, $\gamma_{b}=\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ is the largest value for $\gamma$ such that the state evolution 3.2 has a fixed point ($\bm{M},\bm{Q}$) with
$\bm{M}={\bm{0}},\quad\bm{Q}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}.$ (32)
The fixed point is stable for $\alpha<\alpha_{c}$ and unstable otherwise.
The derivation of the Lemmas is outlined in Appendix B. This two results indicate a change of behavior of the operator norm of ${\underline{\bm{L}}}$ at the critical value $\alpha_{c}$, leading to the following conjecture.
@@
In the high-dimensional limit $n,d\to\infty$, $\nicefrac{{n}}{{d}}\to\alpha$, the empirical spectral distribution associated to the $pn$ eigenvalues of ${\underline{\bm{L}}}$ converges weakly almost surely to a density whose support is strictly contained in a disk of radius $\gamma_{b}=\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ centered at the origin. Moreover
  • for $\alpha<\alpha_{c}$, $\|{\underline{\bm{L}}}\|_{\mathrm{op}}\xrightarrow[n\to\infty]{a.s.}\gamma_{b}$ and the associated eigenvector is not correlated with $\bm{W}_{\star}$;
  • for $\alpha>\alpha_{c}$, $\|{\underline{\bm{L}}}\|_{\mathrm{op}}\xrightarrow[n\to\infty]{a.s.}\gamma_{s}>\gamma_{b}$ and the associated eigenvector defined in (16), weakly recovers the signal.
This conjecture, motivated by the results (3.3-3.4) is perfectly supported by extensive simulations, as illustrated in Fig. 1, 2. A rigorous proof requires, however, a finer control of the spectral norm of these operators.
@@ 3.2 Symmetric spectral method
In order to simplify the notation, define the symmetric $p\times p$ matrix
$\bm{\mathcal{T}}(y)\coloneqq{\bm{I}}-\mathrm{Cov}^{-1}[\bm{\mathbf{z}}\big{|}y].$ (33)
@@
Consider the linear GAMP algorithm (55,56) with denoiser functions
${\bm{g}}_{\rm out}^{t}(\bm{\omega},y)$ $=\bm{\mathcal{T}}(y){\bm{V}}_{t}\left(a_{t}{\bm{I}}-{\bm{V}}_{t}\bm{\mathcal{T}}({\mathrm{y}}){\bm{V}}_{t}\right)^{-1}\bm{\omega},$ (34)
$\bm{f}_{in}^{t+1}(\bm{b})$ $={\bm{V}}_{t+1}\bm{b},$ (35)
@@
$a_{t}$ a parameter to be fixed and $\gamma_{t}$ chosen, a posteriori, such that $\|{\bm{V}}_{t+1}\|_{\mathrm{op}}=1$. Note that ${\bm{V}}_{t}$ is symmetric at all times given a symmetric initialization ${\bm{V}}_{0}$.
In Appendix C we show that, for properly chosen $a_{t}$, the fixed point for the iterate $\bm{\mathbf{\hat{W}}}^{t}{\bm{V}}_{t}$ is the eigenvector of ${\underline{\bm{T}}}$ with eigenvalue $\lim_{t\to\infty}a_{t}\gamma_{t}$. The denoiser functions chosen for this GAMP algorithm are derived as a generalization of the ones in Zhang et al. (2024), where a similar approach has been used to characterize the recovery properties of spectral algorithms for structured single-index models.
Proposition 3.7 (State evolution Javanmard and Montanari (2013); Mondelli and Montanari (2018)).
@@
$\bm{M}^{t+1}$ $=\alpha{\mathcal{F}}({\bm{M}}^{t};a_{t},{\bm{V}}_{t}),$ (37)
$\bm{Q}^{t+1}$ $={\bm{M}}^{t}({\bm{M}}^{t})^{T}+\alpha({\mathcal{G}}({\bm{M}}^{t};a_{t},{\bm{V}}_{t})+\tilde{{\mathcal{F}}}({\bm{Q}}^{t};a_{t},{\bm{V}}_{t})),$
@@
while $\tilde{{\mathcal{F}}}$ and ${\mathcal{G}}$ are given in eq. (97,98).
@@
Lemma 3.8.
For $\alpha>\alpha_{c}$, consider $a_{t}=a>1$ solution of $\nu_{1}^{{\mathcal{F}}}(a_{t})=\alpha^{-1}$ and $\gamma_{t}$ such that $\|{\bm{V}}_{t+1}\|_{\mathrm{op}}=1$, with ${\bm{V}}_{t+1}$ given by eq. (36). Then, the state evolution has a stable fixed point $({\bm{M}},{\bm{Q}})$ with
${\bm{M}}\neq{\bm{0}},\quad{\bm{Q}}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\},$ (39)
correspondent to the eigenvalue $\lambda_{s}=a\gamma$.
Lemma 3.9.
For $\alpha\geq\alpha_{c}$, $a_{t}=a\geq 1$ solution to $\nu_{1}^{\tilde{{\mathcal{F}}}}(a)=\alpha^{-1}$, $\gamma_{t}$ such that $\|{\bm{V}}_{t+1}\|_{\mathrm{op}}=1$, with ${\bm{V}}_{t+1}$ given by eq. (36), the state evolution has an unstable fixed point $({\bm{M}},{\bm{Q}})$ with
${\bm{M}}={\bm{0}},\quad{\bm{Q}}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\},$ (40)
correspondent to the eigenvalue $\lambda_{b}=a\gamma$.
The derivation of the Lemmas is outlined in Appendix C.
@@
In the high-dimensional limit $n,d\to\infty$, $\nicefrac{{n}}{{d}}\to\alpha$, for $\alpha>\alpha_{c}$, the largest eigenvalue of ${\underline{\bm{T}}}$ converges to $\lambda_{s}$, defined in Lemma 3.8. In this regime, the symmetric spectral method, defined in (17), weakly recovers the signal. Moreover, the empirical spectral distribution of the $pd$ eigenvalues of ${\underline{\bm{T}}}$ converges weakly almost surely to a density upper bounded by $\lambda_{b}<\lambda_{s}$ defined in Lemma 3.9.
By construction, in multi-index models ($p>1$), the tensor ${\underline{\bm{T}}}$, or equivalently its flattened matrix representation $\bm{flat}({\underline{\bm{T}}})$, exhibits a highly structured form due to the presence of repeated entries from the measurement matrix $\bm{\mathbf{X}}$. This intrinsic redundancy complicates its analysis using standard random matrix theory tools. Conjecture 3.10 based on results (3.8-3.9) and further supported by numerical simulations (see Fig. 2, 3), offers a novel framework for understanding the spectral properties of such matrices. Moreover, for any link function such that $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ admits a common basis for all $y$, with real eigenvalues $\{\lambda_{k}(y)+1\}_{k=1}^{p}$, the analysis of spectrum of $\bm{flat}({\underline{\bm{T}}})$ can be simplified following the arguments in Appendix D.2. Indeed, the symmetric spectral method reduces to the diagonalization of $p$ matrices
@@
Their structure allows the use of the techniques in Mondelli and Montanari (2018); Lu and Li (2019) to analyze the spectrum, and could support the formalization of the results in Conjecture 3.10 for this subset of problems.
Figure 2: Largest eigenvalues (magnitude, if complex) of the tensors ${\underline{\bm{L}}}$ (triangles), $n=5000$, and ${\underline{\bm{T}}}$ (circles), $d=5000$, for the function $g(z_{1},z_{2})=z_{1}z_{2}$, versus the sample complexity $\alpha$. The orange and blue lines, respectively represent the values of the largest eigenvector $\nicefrac{{\alpha}}{{\alpha_{c}}}$ and the edge of the bulk $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ in Conj. 3.5 for the asymmetric spectral method. The green and purple line correspond to the values of the largest eigenvalue $\lambda_{s}$ and the edge of the bulk $\lambda_{b}$ in Conj. 3.10 for the symmetric spectral method.
Figure 3: Distribution of the eigenvalues of ${\underline{\bm{T}}}$, $d=10^{4}$, for the link function $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$, $p=4$. The critical threshold in $\alpha_{c}=2$. The distribution is truncated on the left. (Left) $\alpha=1<\alpha_{c}$. (Center) $\alpha=\alpha_{c}$. (Right) $\alpha=6>\alpha_{c}$. As predicted by the state evolution framework, in this regime we observe four eigenvalues (highlighted in green) separated from the main bulk, centered around $\lambda_{s}$ (green vertical line) obtained in Lemma 3.8. The vertical purple line correspond to the value $\lambda_{b}$ provided in Lemma 3.9 as a bound for the bulk.

### v2

@@ 2 Setting and Notations
$\delta\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\delta\bm{\mathbf{\hat{W}}}^{t}-{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\delta\bm{\Omega}^{t-1}\right)\right),$ (14)
$\delta\bm{\mathbf{\hat{W}}}^{t+1}$ $=\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\delta\bm{\Omega}^{t}\right)\right)$ (15)
Substituting the second equation into the first, this is equivalent to ${\rm vec}\left(\delta\bm{\Omega}^{t+1}\right)=\bm{L}\,{\rm vec}\left(\delta\bm{\Omega}^{t}\right)$. Moreover, assuming convergence, [^1] Here and in the rest of the paper, we drop the time index $t$ to refer to the quantities at convergence. one ends up with ${\rm vec}\left(\delta\bm{\mathbf{\hat{W}}}\right)=\bm{T}\,{\rm vec}\left(\delta\bm{\mathbf{\hat{W}}}\right)$.
This suggests that at first leading order in the estimates, the dynamics is governed by power-iteration (and respectively a variant of it) on the matrix $\bm{L}$ (respectively the symmetric matrix $\bm{T}$) defined previously. As power iteration converges under mild assumptions to the top (matrix) eigenvector of the tensor, this further suggests to use the top eigenvectors $\bm{\mathbf{\hat{W}}}_{\bm{L}}$, $\bm{\mathbf{\hat{W}}}_{\bm{T}}$ of the corresponding matrices as an estimate for the weight matrix $\bm{W}_{\star}$.
Additional details on the linearized GAMP are reported in Appendix A.1. Similar approaches have been thoroughly investigated in the context of single-index models (Mondelli and Montanari, 2018; Maillard et al., 2022), community detection (Krzakala et al., 2013; Saade et al., 2014), spiked matrix estimation (Lesieur et al., 2017; Aubin et al., 2019), where they have been provably shown to provide a non-vanishing correlation with the ground truth exactly at the optimal weak recovery threshold. It is interesting to notice that the spectral estimators $\bm{\mathbf{\hat{W}}}_{\bm{L}}$ and $\bm{\mathbf{\hat{W}}}_{\bm{T}}$ correspond to the generalization for multi-index models of the spectral methods derived in (Maillard et al., 2022), respectively from the linearization of the optimal Vector Approximate Message Passing (Schniter et al., 2016; Rangan et al., 2017) and the Hessian of the TAP free energy associated to the posterior distribution for the weights (Saade et al., 2014). In particular, for $p\!=\!1$, the matrices proposed in our manuscript exactly reduce to the two ones (called respectively "TAP" and "LAMP") investigated in (Maillard et al., 2022).
@@ 3 Main Technical Results
In order to characterize the weak recovery of the proposed spectral methods, we define two message passing schemes tailored to respectively have eigenvectors of $\bm{L}$ and $\bm{T}$ as fixed points. Similarly to these previous works, we leverage the state evolution associated to the algorithms in order to quantify the alignment between the spectral estimators and the weight matrix $\bm{W}_{\star}$, tracking the overlap matrices
@@ 3.1 Asymmetric spectral method
For $\gamma>0$, consider the linear GAMP algorithm (37,38)
$\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-\gamma^{-1}{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\bm{\Omega}^{t-1}\right)\right),$ (17)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\gamma^{-1}\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\bm{\Omega}^{t}\right)\right).$ (18)
Figure 1: Distribution of the eigenvalues (dots) $\lambda\in\mathbb{C}$ of $\bm{L}$ at finite $n=10^{4}$, for $g(z_{1},z_{2})=z_{1}z_{2}$, $\alpha_{c}\approx 0.59375$. (Left) $\alpha=0.4<\alpha_{c}$. (Right) $\alpha=1>\alpha_{c}$. The dashed blue circle has radius equal to $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$, i.e. the value $\gamma_{b}$ predicted in Theorem 3.4. The dashed orange vertical line corresponds to $\operatorname{Re}\lambda=\nicefrac{{\alpha}}{{\alpha_{c}}}$, the eigenvalue $\gamma_{s}$ defined in Theorem 3.3. As predicted by the state evolution equations for this problem, two significant eigenvalues (highlighted in orange) are observed near this vertical line.
When $\gamma$ is chosen as an eigenvalue of $\bm{L}$, the correspondent eigenvector $\bm{\omega}\in\mathbb{R}^{np}$ is a fixed point of eq. (17) for $\bm{\Omega}^{t}={\rm mat}\left(\bm{\omega}\right)$. In the high-dimensional limit, the asymptotic overlaps of the estimator $\bm{\mathbf{\hat{W}}}$ can be tracked thanks to the state evolution equations, which follow from an immediate application of the general result of Javanmard and Montanari (2013) to the GAMP algorithm (3.6).
Proposition 3.2 (State evolution (Javanmard and Montanari, 2013)).
@@
$\bm{M}^{t+1}$ $=\alpha\gamma^{-1}{\mathcal{F}}(\bm{M}^{t});$ (19)
@@
${\mathcal{F}}(\bm{M})\coloneqq\mathbb{E}_{{\mathrm{y}}}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})],\qquad{\mathcal{G}}(\bm{M})\coloneqq\mathbb{E}_{{\mathrm{y}}}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})\bm{M}^{T}{\bm{G}}({\mathrm{y}})].$ (21)
With the state evolution equations in hand, one can derive a sharp characterization of the asymptotic weak recovery threshold in terms of the spectral properties of the estimator by a linear stability argument (Strogatz, 2001)
Theorem 3.3.
For $\alpha>\alpha_{c}$, $\gamma_{s}=\nicefrac{{\alpha}}{{\alpha_{c}}}$ is the largest value of $\gamma$ such that the state evolution 3.2 has a stable fixed point ($\bm{M},\bm{Q}$) with $\bm{M}\neq{\bm{0}},\;\bm{Q}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$. Additionally, $\bm{M}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$.
Theorem 3.4.
For all $\alpha\in\mathbb{R}$, $\gamma_{b}=\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ is the largest value for $\gamma$ such that the state evolution 3.2 has a fixed point ($\bm{M},\bm{Q}$) with $\bm{M}={\bm{0}},\;\bm{Q}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$. The fixed point is stable for $\alpha<\alpha_{c}$ and unstable otherwise.
The derivation of the Theorems is outlined in Appendix B, where we further show that the iterations of Algorithm 3.1 correspond to the power-iteration on the operator $\bm{L}$, normalized by $\gamma$. Assuming the convergence of Algorithm 3.1 in $\mathcal{O}(\log d)$ iterations, Theorem 3.3 thus implies that for $\alpha>\alpha_{c}$ the top-most eigenvector of $\bm{L}$ achieves a non-vanishing overlap along $\bm{W}_{\star}$, with the eigenvalues converging to $\gamma_{s}$. Analogously, Theorem 3.4 implies that for $\alpha<\alpha_{c}$, the top-most eigenvector of $\bm{L}$ has a vanishing overlap along $\bm{W}_{\star}$, with the eigenvalues converging to $\gamma_{b}$.
Thus, the above two results indicate a change of behavior of the operator norm of $\bm{L}$ at the critical value $\alpha_{c}$, leading to the following conjecture (a detailed sketch can be found in Appendix D):
@@
In the high-dimensional limit $n,d\to\infty$, $\nicefrac{{n}}{{d}}\to\alpha$, the empirical spectral distribution associated to the $pn$ eigenvalues of $\bm{L}$ converges weakly almost surely to a density whose support is strictly contained in a disk of radius $\gamma_{b}=\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ centered at the origin. Moreover
  • for $\alpha<\alpha_{c}$, $\|\bm{L}\|_{\mathrm{op}}\xrightarrow[n\to\infty]{a.s.}\gamma_{b}$ and the associated eigenvector is not correlated with $\bm{W}_{\star}$;
  • for $\alpha>\alpha_{c}$, $\|\bm{L}\|_{\mathrm{op}}\xrightarrow[n\to\infty]{a.s.}\gamma_{s}>\gamma_{b}$ and the associated eigenvector defined in (10), weakly recovers the signal.
This conjecture, motivated by the results (3.3-3.4) is perfectly supported by extensive simulations, as illustrated in Fig. 1, 2. An entire rigorous proof requires, however, a fine control of the spectral norm of these operators, which is a notably difficult problem in random matrix theory.
@@ 3.2 Symmetric spectral method
In order to simplify the notation, define the symmetric matrices $\bm{\mathcal{T}}(y)\in\mathbb{R}^{p\times p}$ and $\hat{{\bm{G}}}_{\bm{\mathcal{T}}}^{t}\in\mathbb{R}^{np\times np}$
$\bm{\mathcal{T}}(y)\coloneqq{\bm{G}}(y)\left({\bm{G}}(y)+{\bm{I}}\right)^{-1}={\bm{I}}-\mathrm{Cov}^{-1}[\bm{\mathbf{z}}\big{|}y].$ (22)
$\left[\hat{{\bm{G}}}_{\bm{\mathcal{T}}}^{t}\right]_{(i\mu),(j\nu)}\coloneqq\delta_{ij}[\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}_{t}\left(a_{t}{\bm{I}}-{\bm{V}}_{t}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}_{t}\right)^{-1}]_{\mu\nu},\quad i,j\in\llbracket n\rrbracket,\,\mu\nu\in\llbracket p\rrbracket.$ (23)
@@
Consider the linear GAMP algorithm (37,38)
$\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-{\rm mat}\left(\hat{{\bm{G}}}_{\bm{\mathcal{T}}}^{t-1}\,{\rm vec}\left(\bm{\Omega}^{t-1}\right)\right){\bm{V}}_{t},$ (24)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\left(\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{{\bm{G}}}_{\bm{\mathcal{T}}}^{t}\,{\rm vec}\left(\bm{\Omega}^{t}\right)\right)-\bm{\mathbf{\hat{W}}}^{t}{\bm{A}}_{t}\right){\bm{V}}_{t+1}$ (25)
@@
$[{\bm{A}}_{t}]_{\mu\nu}$ $\!=\!d^{-1}\sum_{i\in\llbracket n\rrbracket}\left[\hat{{\bm{G}}}_{\bm{\mathcal{T}}}\right]_{(i\mu),(i\nu)},\quad\mu,\nu\in\llbracket p\rrbracket,$ (27)
$a_{t}$ a parameter to be fixed and $\gamma_{t}$ chosen, a posteriori, such that $\|{\bm{V}}\|_{\mathrm{op}}=1$. Note that ${\bm{V}}_{t}$ is symmetric at all times given a symmetric initialization ${\bm{V}}_{0}$.
In Appendix C we show that, for properly chosen $a_{t}$, the fixed point for the iterate $\bm{\mathbf{\hat{W}}}^{t}{\bm{V}}_{t}$ is the eigenvector of $\bm{T}$ with eigenvalue $\lim_{t\to\infty}a_{t}\gamma_{t}$. The denoiser functions chosen for this GAMP algorithm are derived as a generalization of the ones in (Zhang et al., 2024), where a similar approach has been used to characterize the recovery properties of spectral algorithms for structured single-index models.
Proposition 3.7 (State evolution (Javanmard and Montanari, 2013; Mondelli and Montanari, 2018)).
@@
$\bm{M}^{t+1}=\alpha{\mathcal{F}}({\bm{M}}^{t};a_{t},{\bm{V}}_{t}),\qquad\bm{Q}^{t+1}={\bm{M}}^{t}({\bm{M}}^{t})^{T}+\alpha({\mathcal{G}}({\bm{M}}^{t};a_{t},{\bm{V}}_{t})+\tilde{{\mathcal{F}}}({\bm{Q}}^{t};a_{t},{\bm{V}}_{t})),$ (28)
@@
while $\tilde{{\mathcal{F}}}$ and ${\mathcal{G}}$ are given in eq. (81,82).
@@
Theorem 3.8.
For $\alpha>\alpha_{c}$, consider $a_{t}=a>1$ solution of $\nu_{1}^{{\mathcal{F}}}(a_{t})=\alpha^{-1}$ and $\gamma_{t}$ such that $\|{\bm{V}}\|_{\mathrm{op}}=1$, with ${\bm{V}}_{t+1}$ given by eq. (26). Then, the state evolution has a stable fixed point $({\bm{M}},{\bm{Q}})$ with ${\bm{M}}\neq{\bm{0}},\;{\bm{Q}}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$, correspondent to the eigenvalue $\lambda_{s}=a\gamma$.
Theorem 3.9.
For $\alpha\geq\alpha_{c}$, $a_{t}=a\geq 1$ solution to $\nu_{1}^{\tilde{{\mathcal{F}}}}(a)=\alpha^{-1}$, $\gamma_{t}$ such that $\|{\bm{V}}\|_{\mathrm{op}}=1$, with ${\bm{V}}_{t+1}$ given by eq. (26), the state evolution has an unstable fixed point $({\bm{M}},{\bm{Q}})$ with ${\bm{M}}={\bm{0}},\;{\bm{Q}}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$, corresponding to the eigenvalue $\lambda_{b}=a\gamma$.
The derivation of the Theorems is outlined in Appendix C. Analogously to how Theorems 3.3, 3.4 motivate Conjecture 3.5, we show in Appendix C that Theorems 3.8, 3.9, along with a mapping to power-iteration, lead to the following conjecture:
@@
In the high-dimensional limit $n,d\to\infty$, $\nicefrac{{n}}{{d}}\to\alpha$, for $\alpha>\alpha_{c}$, the largest eigenvalue of $\bm{T}$ converges to $\lambda_{s}$, defined in Theorem 3.8. In this regime, the symmetric spectral method, defined in (12), weakly recovers the signal. Moreover, the empirical spectral distribution of the $pd$ eigenvalues of $\bm{T}$ converges weakly almost surely to a density upper bounded by $\lambda_{b}<\lambda_{s}$ defined in Theorem 3.9.
We emphasize that, when $p=1$, the proposed symmetric estimator specializes to the method in Mondelli and Montanari (2018); Lu and Li (2019); Barbier et al. (2019) for the single-index model, and therefore benefits from a fully rigorous characterization of its weak recovery properties and spectral phase transitions. By construction, in multi-index models ($p>1$), the matrix $\bm{T}$ exhibits a highly structured form due to the presence of repeated entries from the measurement matrix $\bm{\mathbf{X}}$. This intrinsic redundancy complicates its analysis using standard random matrix theory tools. Conjecture 3.10 based on results (3.8-3.9) and further supported by numerical simulations (see Fig. 2, 3), offers a novel framework for understanding the spectral properties of such matrices.
Moreover, for any model $\mathsf{P}(\cdot|\bm{\mathbf{z}})$ such that ${\bm{G}}(y)=\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}\big{|}{\mathrm{y}}=y]$ admits a common basis for all $y$, with real eigenvalues $\{\lambda_{k}(y)\}_{k=1}^{p}$, the analysis of spectrum of $\bm{T}$ can be simplified following the arguments in Appendix C.3.1. Indeed, the symmetric spectral method reduces to the diagonalization of $p$ matrices
@@
Their structure allows the use of the techniques in (Mondelli and Montanari, 2018; Lu and Li, 2019) to analyze the spectrum, and supports the formalization of the results in Conjecture 3.10 for this subset of problems. In Appendix C.3 we prove the following result
Theorem 3.11.
Assume that the matrix $\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|y]$ admits a basis of orthonormal eigenvectors independent of $y$. Then, in the high-dimensional limit $n,d\to\infty$, $\nicefrac{{n}}{{d}}\to\alpha$, for $\alpha>\alpha_{c}$, the largest eigenvalue of $\bm{T}$ converges to $\lambda_{s}=1$. Moreover, the empirical spectral distribution of the $pd$ eigenvalues of $\bm{T}$ converges weakly almost surely to a density upper bounded by $\lambda_{b}<1$ defined in Theorem 3.9.
Figure 2: Largest eigenvalues (magnitude, if complex) of the matrices $\bm{L}$ (triangles), $n=5000$, and $\bm{T}$ (circles), $d=5000$, for the function $g(z_{1},z_{2})=z_{1}z_{2}$, versus the sample complexity $\alpha$. The orange and blue lines, respectively represent the values of the largest eigenvector $\nicefrac{{\alpha}}{{\alpha_{c}}}$ and the edge of the bulk $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$ in Conj. 3.5 for the asymmetric spectral method. The green and purple line correspond to the values of the largest eigenvalue $\lambda_{s}$ and the edge of the bulk $\lambda_{b}$ in Conj. 3.10 for the symmetric spectral method.
Figure 3: Distribution of the eigenvalues of $\bm{T}$, $d=10^{4}$, for the link function $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$, $p=4$. The critical threshold in $\alpha_{c}=2$. The distribution is truncated on the left. (Left) $\alpha=1<\alpha_{c}$. (Center) $\alpha=\alpha_{c}$. (Right) $\alpha=6>\alpha_{c}$. As predicted by the state evolution, we observe four eigenvalues (in green) separated from the main bulk, centered around $\lambda_{s}=1$ (green vertical line) obtained in Theorem 3.8. The vertical purple line correspond to the value $\lambda_{b}$ provided in Theorem 3.9 as a bound for the bulk.